solve this exercise using the fact that the sum of the measures of the three angles of a triangle is 180 degrees. In a triangle, the measures of the three angles are x, x-29, and x+35. what is the measure of each angle?

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39630)

Make the sum of those expressions equal to 180 and just solve the equation!

Answer by ikleyn(52914)

.
solve this exercise using the fact that the sum of the measures of the three angles of a triangle is 180 degrees.
In a triangle, the measures of the three angles are x, x-29, and x+35. what is the measure of each angle?
~~~~~~~~~~~~~~~~~~~~~~

Write equation as instructed

    x + (x-29) + (x+35) = 180.


    +-----------------------------------------+
    |      For it, you should not think.      |
    |    You simply should do as instructed.  |
    +-----------------------------------------+


Now you should solve this equation.


For it, you rewrite it, moving the constant terms from the left side to the right side, changing the signs to opposite

    x + x + x = 180 + 29 - 35

or

       3x     =      174

        x     =      174/3 = 58.


So, one angle is 58 degrees;  second angle is 58-29 = 29 degrees and third angle is 58+35 = 93 degrees.


ANSWER.  The angles are  58 degrees; 29 degrees and 93 degrees.

Solved.

-----------------

This problem is EXTREMELY simple.

What difficult did you find in it, that forced you to write to the forum and ask for help ?